# Proton Lifetime in Minimal Supersymmetric SU(5) with Gauge Mediation

### v1

@@ II.1 The Grand Unified Theory
We consider the following $SU(5)$ invariant superpotential:
@@ II.2 Messenger sector
where $\Psi_{L}$ and $\Psi_{D}$ are the $SU(2)_{L}$ doublet and $SU(3)_{C}$ triplet messengers, and $Z$ is a singlet superfield which breaks SUSY [^4] Although GUT theories dictate that $M_{L}$ and $M_{D}$ arise from the same mass term, $M\Psi\bar{\Psi}=M(\Psi_{D}\Psi_{\bar{D}}+\Psi_{L}\Psi_{\bar{L}})$, we can achieve $M_{L}\neq M_{D}$ by introducing additional terms of the form $\lambda_{\Psi}\bar{\Psi}\Sigma\Psi$. Once $\Sigma$ gets a VEV, the masses are split, $M_{L}=M-3\lambda_{\Psi}V$ and $M_{D}=M+2\lambda_{\Psi}V$.. When integrated out, these messengers generate the following gaugino and scalar masses
@@
where $C_{2,3}(i)$ are the quadratic Casimir invariants of the field $i$ for $SU(2)_{L}$ and $SU(3)_{C}$, respectively, $\Lambda_{L,D}\equiv k_{L,D}\langle F_{Z}\rangle/M_{L,D}$ with $\langle F_{Z}\rangle=\xi_{Z}$, and $\Lambda_{Y}^{2}=(2/5)\Lambda_{D}^{2}+(3/5)\Lambda_{L}^{2}$. We assume that the messenger masses are of the same order, i.e. $M_{\rm mess}\sim M_{L}\sim M_{D}$. Note that for the other SUSY breaking parameters, including $A$-terms, there are contributions from the messengers, but they are much smaller than the gaugino masses and so are taken to be zero in our numerical calculation.
@@
where $H_{1,2}$ are two Higgs doublet fields. In order to arrive at a viable model, this problem must be solved. However, in the context of GUT theories with low messenger scales, this problem can be easily solved. When combining GMSB with grand unification there is another problem with $\mu$ and $B_{\mu}$ associated with the GUT matching conditions found in Ellis et al. (2017). First, the $\mu_{H}$ must be tuned against a GUT scale contribution $\lambda V$ to get a weak scale Higgs doublet. This tuning gives us the freedom to choose $\mu$ as we wish. $B_{\mu}$, on the other hand, gets a gravity mediated contribution which is proportional to the gravitino mass, which is small for low scale messenger case. However, the gravity mediated contribution to $B_{\mu}$ is also enhanced by $\mu_{H}$ which must be of order the GUT scale. Generically, this gives $B_{\mu}\sim m_{3/2}M_{GUT}$, which is too larger. This means we can use the same techniques used in gravity mediation found in Ellis et al. (2017) to realize a viable $B_{\mu}$ parameter. Although this requires us to tune $B_{\mu}$, just like what is done in gravity mediated models, the required tuning is significantly less. Thus, GMSB models of $SU(5)$ have some advantages over gravity mediation as well as GMSB without grand unification.
@@ IV Numerical results
In this section, we summarize our numerical results for the proton lifetime, $p\to K^{+}\bar{\nu}$, in the GMSB scenario. We assume that at the messenger scale, the gaugino and scalar masses are given by Eqs. (9)-(11), and $A$-terms and $B$ parameter are subleading and so are set to zero. The RGEs then evolve the parameters to the low-scale [^6] We use the SSARD code for this calculation SSA . The $\mu$ and $B_{\mu}$ parameters are determined by the electroweak symmetry breaking (EWSB) conditions leaving eight free parameters:
@@
It is clear that green and yellow bands for the Higgs mass have a non-trivial dependence on $\Lambda_{D,L}$. The key factor being the stop mass, which can be understood as follows. First, for the $\Lambda_{L}\lesssim 3200$ TeV region, the stop mass is determined solely by $\Lambda_{D}$, and hence, the Higgs mass bands are almost independent of $\Lambda_{L}$. For $3200~{}{\rm TeV}\lesssim\Lambda_{L}\lesssim 4000$ TeV, the $\Lambda_{L}$ contribution to the stop masses and slightly smaller $\Lambda_{D}$ reproduce the needed stop masses. For $4000~{}{\rm TeV}\lesssim\Lambda_{D}$, radiative corrections from stop mixing parameter ($X_{t}=A_{t}-\mu/\tan\beta$) affect the prediction of the Higgs mass. This feature is mainly coming from the value of $\mu$: for $\Lambda_{L}\simeq 4000$ TeV, $\mu=\mathcal{O}(1)$ TeV, and increases with $\Lambda_{L}$. $A_{t}$, on the other hand, is almost completely determined by $\Lambda_{D}$. Hence, $X_{t}$ increases with $\Lambda_{L}$ for fixed $\Lambda_{D}$ in the range of $\Lambda_{L}>4000$ TeV. We numerically check that the radiative corrections associated with $X_{t}$ decreases the SM Higgs mass for larger values of $\Lambda_{L}$. As a result, slightly larger $\Lambda_{D}$ is required to reproduce 125 GeV Higgs mass.

### v2

@@ II.1 The Grand Unified Theory
We consider the following $SU(5)$ invariant superpotential [^4] In our study, we omit other higher dimensional operators assuming that they should contribute at a subleading order. In order for the neglected higher dimensional operators to be important, they would need to dominate the tree-level contribution.:
@@ II.2 Messenger sector
where $\Psi_{L}$ and $\Psi_{D}$ are the $SU(2)_{L}$ doublet and $SU(3)_{C}$ triplet messengers, and $Z$ is a singlet superfield which breaks SUSY [^5] Although GUT theories dictate that $M_{L}$ and $M_{D}$ arise from the same mass term, $M\Psi\bar{\Psi}=M(\Psi_{D}\Psi_{\bar{D}}+\Psi_{L}\Psi_{\bar{L}})$, we can achieve $M_{L}\neq M_{D}$ by introducing additional terms of the form $\lambda_{\Psi}\bar{\Psi}\Sigma\Psi$. Once $\Sigma$ gets a VEV, the masses are split, $M_{L}=M-3\lambda_{\Psi}V$ and $M_{D}=M+2\lambda_{\Psi}V$. By taking $\lambda_{\Psi}\ll 1$, independent masses with $M_{L}\sim M_{D}$ can be realized at any scale.. When integrated out, these messengers generate the following gaugino and scalar masses
@@
where $C_{2,3}(i)$ are the quadratic Casimir invariants of the field $i$ for $SU(2)_{L}$ and $SU(3)_{C}$, respectively, $\Lambda_{L,D}\equiv k_{L,D}\langle F_{Z}\rangle/M_{L,D}$ with $\langle F_{Z}\rangle=\xi_{Z}$, and $\Lambda_{Y}^{2}=(2/5)\Lambda_{D}^{2}+(3/5)\Lambda_{L}^{2}$. We assume that the messenger masses are of the same order, i.e. $M_{\rm mess}\sim M_{L}\sim M_{D}$ [^6] These masses are corrected by the scalar vev of the supersymmetry breaking field, $\langle\phi_{Z}\rangle$, $M_{L,D}+k_{L,D}\langle\phi_{Z}\rangle$. However, the scalar vev of $Z$ is expected to be of order the SUSY breaking scale. In our calculation, we take $M_{\rm mess}=\mathcal{O}(10^{3})$ TeV, while our SUSY breaking scale is $\mathcal{O}(10)$ TeV. Thus, our assumption of $M_{L,D}\gg k_{L,D}\langle\phi_{Z}\rangle$ is justified.. Note that for the other SUSY breaking parameters, including $A$-terms, there are contributions from the messengers, but they are much smaller than the gaugino masses and so are taken to be zero in our numerical calculation.
@@
where $H_{1,2}$ are two Higgs doublet fields. In order to arrive at a viable model, this problem must be solved. However, in the context of GUT theories with low messenger scales, this problem can be easily solved. When combining GMSB with grand unification there is another problem with $\mu$ and $B_{\mu}$ associated with the GUT matching conditions found in Ellis et al. (2017). First, the $\mu_{H}$ must be tuned against a GUT scale contribution $\lambda V$ to get a weak scale Higgs doublet. This tuning gives us the freedom to choose $\mu$ as we wish. $B_{\mu}$, on the other hand, gets a gravity mediated contribution which is proportional to the gravitino mass, which is small for low scale messenger case. However, the gravity mediated contribution to $B_{\mu}$ is also enhanced by $\mu_{H}$ which must be of order the GUT scale. Generically, this gives $B_{\mu}\sim m_{3/2}M_{GUT}$, which is too large. This means we can use the same techniques used in gravity mediation found in Ellis et al. (2017) to realize a viable $B_{\mu}$ parameter. Although this requires us to tune $B_{\mu}$, just like what is done in gravity mediated models, the required tuning is significantly less. Thus, GMSB models of $SU(5)$ have some advantages over gravity mediation as well as GMSB without grand unification.
@@ IV Numerical results
[…]
@@
It is clear that green and yellow bands for the Higgs mass have a non-trivial dependence on $\Lambda_{D,L}$. The key factor being the stop mass, which can be understood as follows. First, for the $\Lambda_{L}\lesssim 3200$ TeV region, the stop mass is determined solely by $\Lambda_{D}$, and hence, the Higgs mass bands are almost independent of $\Lambda_{L}$. For $3200~{}{\rm TeV}\lesssim\Lambda_{L}\lesssim 4000$ TeV, the $\Lambda_{L}$ contribution to the stop masses and slightly smaller $\Lambda_{D}$ reproduce the needed stop masses. For $4000~{}{\rm TeV}\lesssim\Lambda_{D}$, radiative corrections from stop mixing parameter ($X_{t}=A_{t}-\mu/\tan\beta$ with $A_{t}\equiv A_{t}(M_{\rm SUSY})$ [^13] Non-zero $A_{t}$ at the SUSY scale can be obtained by the RGE evolution from $M_{\rm mess}$ to $M_{\rm SUSY}$, even when we set all $A$-terms at $M_{\rm mess}$ to be zero.) affect the prediction of the Higgs mass. This feature is mainly coming from the value of $\mu$: for $\Lambda_{L}\simeq 4000$ TeV, $\mu=\mathcal{O}(1)$ TeV, and increases with $\Lambda_{L}$. $A_{t}$, on the other hand, is almost completely determined by $\Lambda_{D}$. Hence, $X_{t}$ increases with $\Lambda_{L}$ for fixed $\Lambda_{D}$ in the range of $\Lambda_{L}>4000$ TeV. We numerically check that the radiative corrections associated with $X_{t}$ decreases the SM Higgs mass for larger values of $\Lambda_{L}$. As a result, slightly larger $\Lambda_{D}$ is required to reproduce 125 GeV Higgs mass.